# Evidence that Mass Loss on the Red Giant Branch Decreases with Metallicity

### v1

@@ Abstract
Mass loss on the red giant branch (RGB) influences stellar evolution, properties of stellar populations, and Galactic chemical enrichment, yet remains poorly constrained observationally. Current models provide limited insight into how stellar properties, particularly how metallicity and mass, affects RGB mass loss. Here, I introduce a new observational approach that uses the age–velocity-dispersion relation and the lower-mass boundary of red giants as precise evolutionary markers. These markers, informed by Galactic evolution, allow us to construct observational isochrones for field stars. By comparing masses of RGB stars and red clump (RC) stars at the same age in the Kepler sample, I derive empirical measurements of integrated RGB mass loss at several points in age and metallicity. Combining these new observational measurements with open-cluster studies, I showed that the integrated mass loss on the RGB decreases with metallicity, and may also decrease with stellar mass. The average mass loss rate, which accounts for RGB lifetimes and the initial mass differences between RGB and RC stars at the same age, also show a similar trend. These findings challenge current mass-loss prescriptions widely adopted in stellar evolutionary models, since none of them is able to produce the observed mass-loss trend without widely adjusting free parameters. This highlights an urgent need to revise mechanisms that govern RGB mass loss.
Asteroseismology (73), Stellar oscillations (1617), Stellar mass loss (1613), Red giant branch (1368), Red giant clump (1370), Milky Way evolution (1052), Open star clusters (1160), Milky Way dynamics (1051), Milky Way Galaxy (1054)
{CJK}
UTF8gbsn
@@ 1 Introduction
Mass loss during the red giant branch (RGB) phase is a critical but still poorly understood aspect of stellar evolution (M. Catelan, 2009). Low- and intermediate-mass stars ($M\leq 2$ $\text{M}_{\odot}$) typically lose up to 0.2 $\text{M}_{\odot}$ by the time they reach the RGB tip (M. H. Pinsonneault et al., 2025). This process has broad implications on inferred stellar ages for red clump (RC) stars (L. Girardi, 2016), horizontal branch morphologies in globular clusters (A. V. Sweigart, 1997; M. Catelan, 2000), and pulsation properties of RR Lyrae stars (V. Caloi & F. D’Antona, 2008; L. Molnár et al., 2024). Mass loss also alters planetary system demographics and dynamics around evolved stars (D. Veras et al., 2013; E. Villaver et al., 2014). In close binaries, it shapes binary interactions at the RGB tip, affecting the formation of extreme horizontal-branch hot subdwarf-B stars (Z. Han et al., 2003).
@@
Unlike the AGB mass loss — where pulsation-driven, dust-enhanced winds are established to be responsible — RGB mass-loss mechanisms remain elusive, thus the descriptions are largely empirical. Reimers’ law, for example, derives mass-loss rate by assuming a fixed fraction of stellar luminosity provides the lost gravitational potential (D. Reimers, 1975). While widely adopted, this relation does not directly describe the underlying physical processes responsible for the mass loss. Mass loss mechanisms generally include pulsation-driven winds, radiation pressure on dust and molecules, magnetic activity or Alfvén-wave driven winds, and binary-induced mass loss. Some prescriptions for the RGB incorporate Alfvén-wave physics (K. P. Schröder & M. Cuntz, 2005; S. R. Cranmer & S. H. Saar, 2011), but these have not been extensively tested in solar-metallicity or thin-disc populations due to a lack of identification for stars of same age. Studying a group of RGB and RC stars at the same age is necessary to characterize integrated RGB mass loss and, so far, such measurements have only been possible for the open clusters observed by Kepler.
In this study, I intend to address this critical gap by proposing new tools to identify field stars that are at the same age, which is informed by Galactic evolution. These tools include the age–velocity-dispersion relation (§2) and the lower mass boundary of red giants (§3). These methods offer extra constraints on RGB mass loss compared to current open cluster constraints (§4). When combined, they yield clear evidence that the integrated mass loss depends on metallicity and mass (§5). In addition, I estimate the average mass-loss rates with stellar models (§6). Based on these constraints, I test common mass-loss prescriptions (§7), discuss the implication for RGB mass-loss mechanisms and finally provide conclusions and outlook (§8).
@@
2 Age-velocity-dispersion relation (AVR)
@@ 2.1 Principle
It has been established that older stars in the Galaxy generally exhibit larger velocity dispersions, while younger stars display smaller dispersions, as a result from dynamical heating over time. This kinematic velocity trend with stellar ages are known as age-velocity-dispersion relation (AVR) (G. Strömberg 1946; R. Wielen 1977; see also J. Bland-Hawthorn & O. Gerhard 2016 for a review).
@@ 2.2 Data and Results
Note. — (1) AVR stands for age-velocity-dispersion relation and LMB stands for lower mass boundary. (2) The values for M67 are adopted from C. Reyes et al. (2025); other values are from this work. (3) Mass refer to the mass at the lower RGB.
@@ 3.1 Principle
Red giants span a broad mass range, but there is typically a lower-mass boundary, below which few are observed. This is because a star must have sufficient mass to evolve off the main sequence within the age of the Galaxy. For a given stellar population (e.g. a Galactic disc with a specified metallicity range), this lower mass boundary effectively marks a fixed age, corresponding to the onset of the formation of that population.
K. Brogaard et al. (2024) applied a similar method to thick disc stars, using the mean mass rather than the lower mass boundary. This can be considered as a valid approach given the approximate single-age (or, equivalently, single-mass) nature of the thick disc stars. For example, A. Miglio et al. (2021) reported a mean age of approximately 11 Gyr with a dispersion of about 1 Gyr for thick disc stars. See also C. Marasco et al. (2025) for a related application to metal-poor stars. They also estimated the integrated mass loss $\Delta M$ by comparing the mean masses of RGB and RC stars.
In more general cases — particularly in the thin disc — assuming a single-age population is not valid. Therefore, focusing on the lower mass boundary, rather than on the mean mass, provides a more robust measure in such populations.
@@ 3.2 Method and Results
The parameter $M_{0}$ is interpreted as the lower mass boundary. Poisson statistics were assumed for the number counts in each bin, resulting in the following likelihood function:
@@ 4 Open clusters
5 Integrated RGB Mass Loss
@@ 5.3 Accuracy of Asteroseismic Scaling Relations
Because Gaia provides independent radius measurements, $\delta(\Delta R)/R$ can be estimated by $(R_{\rm seis}-R_{\rm Gaia})/R_{\rm Gaia}$, permitting a test of this hypothesis. I used $R_{\rm Gaia}$ values from the APOKASC-3 catalog, and computed the difference in this quantity between RGB and RC stars across metallicity bins, as shown in Figure 8. The observed differences are at most 1%, well below the 5% required threshold, thereby ruling out this explanation. This conclusion assumes that the accuracy of the ${\Delta\nu}$ scaling relation is the sole factor contributing to the RGB-RC differences. I tested two other mass scales using different $f_{\Delta\nu}$ corrections reported by M. H. Pinsonneault et al. (2025). Both tests produced values consistent with those shown in Figure 8.
@@
Figure 9: From top to bottom: RGB lifetime $\tau_{\rm RGB}$, initial mass difference $\Delta M_{\rm init}$ and maximum luminosity reached on the RGB, shown as functions of stellar mass and metallicity. Results are presented for three evolutionary models (left to right: MIST, PARSEC, and YY).
@@ 6.1 Stellar Models
I used stellar models to estimate the RGB lifetime ($\tau_{\rm RGB}$), the age–mass gradient ($d\tau/dM$), and the maximum RGB luminosity ($L_{\rm RGB,max}$). I used three commonly-used sets of stellar evolutionary tracks: MIST (rotation version; J. Choi et al., 2016), PARSEC (v1.2; Y. Chen et al., 2015), and YONSEI-YALE (YY; S. Yi et al., 2001; Y.-C. Kim et al., 2002; S. K. Yi et al., 2003; P. Demarque et al., 2004). These models cover metallicities from ${\rm[M/H]}$ = $-1.0$ to $0.5$ and masses from 0.8 to 1.6 $\text{M}_{\odot}$. Using multiple models permits assessment of the robustness of any conclusions, given there are still uncertainties in RGB evolution.
@@
Figure 11: Predicted average mass loss rates from three different prescriptions shown as open symbols. Observed values are shown as filled symbols, following the same conventions as in Figure 7.
Figure 12: Predicted average mass-loss rates from three different prescriptions, shown as colored symbols. The color gradient indicates stellar mass, from light to dark: 1.0, 1.2 and 1.4 $\text{M}_{\odot}$. Observed values are shown as black symbols, following the same conventions as in Figure 7.
@@ 6.2 Trends with Metallicity and Mass
$\begin{split}\langle\dot{M}\rangle/&(10^{-9}M_{\odot}{\rm yr}^{-1})\\ &=(-4.17\pm 0.36)\times{{\rm[M/H]}}+(1.48\pm 0.12).\end{split}$ (18)
7 Testing Mass-loss Prescriptions
@@
Next, I explored the effect of varying the scaling factors — $\eta_{\mathrm{R}}$, $\eta_{\mathrm{SC}}$, and $\eta_{\mathrm{CS}}$ — within each prescription. In this case, I also included models with three different initial masses: 1.0, 1.2, and 1.4 $\text{M}_{\odot}$. As shown in Figure 12, this additional freedom does not resolve the discrepancy. Although all prescriptions produce a mass gradient with a sign consistent with the observed trend with mass, the overall agreement remains poor.

### v2

@@ Abstract
Mass loss on the red giant branch (RGB) influences stellar evolution, properties of stellar populations, and Galactic chemical enrichment, yet remains poorly constrained observationally. Current models provide limited insight into how stellar properties, particularly how metallicity and mass, affect RGB mass loss. Here, I introduce a new observational approach that uses the age-velocity-dispersion relation and the lower-mass boundary of red giants as precise evolutionary markers. These markers, informed by Galactic evolution, allow us to construct observational isochrones for field stars. By comparing masses of RGB stars and red clump (RC) stars at the same age in the Kepler sample, I derive empirical measurements of integrated RGB mass loss at several points in age and metallicity. Combining these new observational measurements with open-cluster studies, I showed that the integrated mass loss on the RGB decreases with metallicity, and may also decrease with stellar mass. The average mass-loss rate, which accounts for RGB lifetimes and the initial mass differences between RGB and RC stars at the same age, also shows a similar trend. These findings challenge current mass-loss prescriptions widely adopted in stellar evolutionary models, since none of them is able to produce the observed mass-loss trend without widely adjusting free parameters. This highlights an urgent need to revise mechanisms that govern RGB mass loss.
Asteroseismology (73), Stellar oscillations (1617), Rapid stellar oscillations (1363); Stellar mass loss (1613), Red giant branch (1368), Red giant clump (1370), Galactic buldge (2041); Milky Way evolution (1052), Open star clusters (1160), Milky Way dynamics (1051), the Milky Way (1054)
@@ 1 Introduction
Mass loss during the red giant branch (RGB) phase is a critical but still poorly understood aspect of stellar evolution (M. Catelan, 2009). Low- and intermediate-mass stars ($\leq 2 M\text{M}_{\odot}$) typically lose up to 0.2 $\text{M}_{\odot}$ by the time they reach the RGB tip (M. H. Pinsonneault et al., 2025). This process has broad implications on inferred stellar ages for red clump (RC) stars (L. Girardi, 2016), horizontal branch morphologies in globular clusters (A. V. Sweigart, 1997; M. Catelan, 2000), and pulsation properties of RR Lyrae stars (V. Caloi & F. D’Antona, 2008; L. Molnár et al., 2024; H. Zhang et al., 2025). Mass loss also alters planetary system demographics and dynamics around evolved stars (D. Veras et al., 2013; E. Villaver et al., 2014). In close binaries, it shapes binary interactions at the RGB tip, affecting the formation of extreme horizontal-branch hot subdwarf-B stars (Z. Han et al., 2003).
@@
Unlike the AGB mass loss — where pulsation-driven, dust-enhanced winds are established to be responsible — RGB mass-loss mechanisms remain elusive, thus the descriptions are largely empirical. Reimers’ law, for example, derives mass-loss rate by assuming a fixed fraction of stellar luminosity provides the lost gravitational potential (D. Reimers, 1975). While widely adopted, this relation does not directly describe the underlying physical processes responsible for the mass loss. Mass loss mechanisms generally include pulsation-driven winds, radiation pressure on dust and molecules, magnetic activity or Alfvén-wave driven winds, and binary-induced mass loss. Some prescriptions for the RGB incorporate Alfvén-wave physics (K. P. Schröder & M. Cuntz, 2005; S. R. Cranmer & S. H. Saar, 2011; T. K. Suzuki et al., 2024), but these have not been extensively tested in solar-metallicity or thin-disc populations due to a lack of identification for stars of same age. Studying a group of RGB and RC stars at the same age is necessary to characterize integrated RGB mass loss and, so far, such measurements have only been possible for the open clusters observed by Kepler.
In this study, I intend to address this critical gap by proposing new tools to identify field stars that are at the same age, which is informed by Galactic evolution. These tools include the age–velocity-dispersion (AVR) relation (§2) and the lower-mass boundary (LMB) of red giants (§3). These methods offer extra constraints on RGB mass loss compared to current open cluster constraints (§4). When combined, they yield clear evidence that the integrated mass loss depends on metallicity and mass (§5). In addition, I estimate the average mass-loss rates with stellar models (§6). Based on these constraints, I test common mass-loss prescriptions (§7), discuss the implication for RGB mass-loss mechanisms and finally provide conclusions and outlook (§8).
@@
2 Age-Velocity-dispersion relation
@@ 2.1 Principle
It has been established that older stars in the Galaxy generally exhibit larger velocity dispersions, while younger stars display smaller dispersions, as a result from dynamical heating over time. This kinematic velocity trend with stellar ages is known as the AVR (G. Strömberg 1946; R. Wielen 1977; see also J. Bland-Hawthorn & O. Gerhard 2016 for a review).
@@ 2.2 Data and Results
Note. — (1) AVR stands for age-velocity-dispersion relation and LMB stands for lower-mass boundary. (2) The values for M67 are adopted from C. Reyes et al. (2025); other values are from this work. (3) Mass refer to the mass at the lower RGB.
@@ 3.1 Principle
Red giants span a broad mass range, but there is typically a lower-mass boundary, below which few are observed. This is because a star must have sufficient mass to evolve off the main sequence within the age of the Galaxy. For a given stellar population (e.g. a Galactic disc with a specified metallicity range), this lower-mass boundary effectively marks a fixed age, corresponding to the onset of the formation of that population.
K. Brogaard et al. (2024) applied a similar method to thick disc stars, using the mean mass rather than the lower-mass boundary. This can be considered as a valid approach given the approximate single-age (or, equivalently, single-mass) nature of the thick disc stars. For example, A. Miglio et al. (2021) reported a mean age of approximately 11 Gyr with a dispersion of about 1 Gyr for thick disc stars. See also C. Marasco et al. (2025) and J. Tayar et al. (2023) for a related application to metal-poor stars. They also estimated the integrated mass loss $\Delta M$ by comparing the mean masses of RGB and RC stars.
In more general cases — particularly in the thin disc — assuming a single-age population is not valid. Therefore, focusing on the lower-mass boundary, rather than on the mean mass, provides a more robust measure in such populations.
@@ 3.2 Method and Results
The parameter $M_{0}$ is interpreted as the lower-mass boundary. Poisson statistics were assumed for the number counts in each bin, resulting in the following likelihood function:
@@ 4 Open Clusters
5 Integrated Red Giant Branch Mass Loss
@@ 5.3 Accuracy of Asteroseismic Scaling Relations
Because Gaia provides independent radius measurements, $\delta(\Delta R)/R$ can be estimated by $(R_{\rm seis}-R_{\rm Gaia})/R_{\rm Gaia}$, permitting a test of this hypothesis. I used $R_{\rm Gaia}$ values from the APOKASC-3 catalog, and computed the difference in this quantity between RGB and RC stars across metallicity bins, as shown in Figure 8. The observed differences are at most 1%, well below the 5% required threshold, thereby ruling out this explanation. This conclusion assumes that the accuracy of the $\Delta\nu$} scaling relation is the sole factor contributing to the RGB-RC differences. I tested two additional mass and radius scales reported by M. H. Pinsonneault et al. (2025), each using a different $f_{\Delta\nu}$ correction. Both radius scales produced results consistent with those shown in Figure 8.
@@
Figure 9: From top to bottom: RGB lifetime $\tau_{\rm RGB}$, initial mass difference $\Delta M_{\rm init}$ and maximum luminosity reached on the RGB, shown as functions of stellar mass and metallicity. Results are presented for three evolutionary models (left to right: MIST, PARSEC, and $Y$).
@@ 6.1 Stellar Models
I used stellar models to estimate the RGB lifetime ($\tau_{\rm RGB}$), the age–mass gradient ($d\tau/dM$), and the maximum RGB luminosity ($L_{\rm RGB,max}$). I used three commonly-used sets of stellar evolutionary tracks: MIST (rotation version; J. Choi et al., 2016), PARSEC (v1.2; Y. Chen et al., 2015), and YONSEI-YALE (YY; S. Yi et al., 2001; Y.-C. Kim et al., 2002; S. K. Yi et al., 2003; P. Demarque et al., 2004). These models cover metallicities from $\rm[M/H]$} = 1.0-1.0 to $0.5$ and masses from 0.8 to 1.6 $\text{M}_{\odot}$. Using multiple models permits assessment of the robustness of any conclusions, given there are still uncertainties in RGB evolution.
@@
Figure 11: Predicted average mass loss rates from three different prescriptions. Observed values are shown as filled symbols, following the same conventions as in Figure 10.
Figure 12: Predicted average mass-loss rates from three different prescriptions, shown as colored symbols. The color gradient represents stellar mass, from light to dark: 1.0, 1.2 and 1.4 $\text{M}_{\odot}$. Observed values are indicated by black symbols, following the same conventions as in Figure 7.
@@ 6.2 Trends with Metallicity and Mass
$\begin{split}\langle\dot{M}\rangle/&(10^{-9}M_{\odot}{\rm yr}^{-1})\\ &=(-4.15\pm 0.36)\times{{\rm[M/H]}}+(1.48\pm 0.12).\end{split}$ (18)
7 Testing the Mass-loss Prescriptions
@@
Next, I explored the effect of varying the scaling factors — $\eta_{\mathrm{R}}$, $\eta_{\mathrm{SC}}$, and $\eta_{\mathrm{CS}}$ — within each prescription. In this case, I also included models with three different initial masses: 1.0, 1.2, and 1.4 $\text{M}_{\odot}$. As shown in Figure 12, this additional freedom does not resolve the discrepancy. Although all prescriptions produce a mass gradient with a sign consistent with the observed trend with mass (A. Theano Theodoridis et al., 2025), the overall agreement remains poor.